A 4.31 kg block of Copper at +275oC is placed into thermal contact with a block of ice at +0.000oC.

When they reach thermal equilibrium, the final temperature of everything is 0.000oC and only HALF of the ice has melted into water!

What was the original amount of ice (in kg)?
Answer #1

Specific heat capacity of copper = 385 J/kg K

Latent heat of fusion of ice = 333.5 J/g

Heat lost by copper = heat gained by ice

mcdT = mL

4.31*385 *( 275-0) = m*333500

m = 1.37 kg

Since only half of the ice has melted

Original amount of ice = 2m = 2*1.37 = 2.74 kg